# Strong-to-weak spontaneous breaking of 1-form symmetry and intrinsically mixed topological order

### v2

@@ IV.4 Path of gapped parent Hamiltonians
To further support the results in the previous subsection, we show below that there is an explicit path of gapped, local Hamiltonians connecting a parent Hamiltonian for $|\rho(\beta_{A})\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}$⟩ to a parent Hamiltonian for $|\rho_{B}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}$⟩. This clearly demonstrates that these two states are in the same phase, and share the same TEE (modulo spurious contributions). The idea is to find an exact unitary operator $\mathcal{U}(\beta_{A})_{m}$ which has the same action on $|\uparrow_{m}\rangle$ as $\prod e^{\beta_{A}A_{v,m}}$ (upon proper renormalization):
$|\rho(\beta_{A})\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}=\mathcal{U}_{m}(\beta_{A})|tc_{u}\rangle|\uparrow_{m}\rangle$ (67)
@@
$H(\beta)$ is a local parent Hamiltonian for $|\rho(\beta_{A})\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}$⟩, whose terms are exponentially localized with a length scale set by $\beta_{A}$.
@@
Since the circuit (LABEL:seqcirc) is only sequential in the $y$ direction only, it leads to a finite correlation length only in the $y$ direction. In other words, $\mathcal{U}_{m}(\beta_{A})O_{e,m}\mathcal{U}_{m}(\beta_{A})^{\dagger}$ generically maps a local operator $O_{e,m}$ supported on site $e$ to an operator with strict locality in the $x$ direction but an exponential tail in the $+y$ direction (note that there is no tail in the $-y$ direction). In Appendix E, we show that $\mathcal{U}(\beta_{A})_{m}$ is locality-preserving for all local operators. In particular, for a verticle link $e$,
@@
is a gapped Hamiltonian that consists of exponentially local terms, as long as $\beta_{A}$ is finite. When $\beta_{A}=0$, the ground state of $H(\beta_{A})$ is simply a toric code fixed point state stacked with a trivial paramagnet, so it clearly has TEE $\log 2$. The same must hold for the ground state $|\rho(\beta_{A})\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}$⟩ of $H(\beta_{A})$ for any finite $\beta_{A}$.
@@ IV.5 Random plaquette term
$|\rho_{\beta}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}\propto\prod_{v}e^{\beta\lambda_{A}(A_{v,u}+A_{v,l})/2}\prod_{p}e^{\beta\lambda_{B}(B_{p,u}+B_{p,l})/2}|\rho_{1}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}},$⟩ , (78)
where $|\rho_{1}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}$⟩ is the Choi state of the identity matrix, which is a state where every spin in $\mathcal{H}_{u}$ is paired with a spin in $\mathcal{H}_{l}$ in a Bell pair. We can write $|\rho_{1}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}$⟩ as the state obtained from full decoherence of toric code with both the $X$ and $Z$ channels.
$|\rho_{1}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}\propto\prod_{e,e^{\prime}}\left(\frac{1+Z_{e,u}Z_{e,l}}{2}\right)\left(\frac{1+X_{e^{\prime},u}X_{e^{\prime},l}}{2}\right)|tc_{u}\rangle|tc_{l}\rangle.$ (79)
As in Sec. IV.3, we will rewrite the state with an additional set of spins acted on by $X_{m},Z_{m}$. Following (61) and (62), $|\rho_{1}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}$⟩ can be written as
$|\rho_{1}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}=|+_{u}\rangle|\uparrow_{m}\rangle$ (80)
Therefore, $|\rho_{\beta}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}$⟩ can be written as
$\begin{split}|\rho_{\beta}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}&\propto\prod_{v}e^{\beta\lambda_{A}(A_{v,u}A_{v,m}+A_{v,m})/2}\\ &\times\prod_{p}e^{\beta\lambda_{B}(B_{p,u}B_{p,m}+B_{p,u})/2}|+_{u}\rangle|\uparrow_{m}\rangle\\ &\propto\left(\prod_{p}e^{\beta\lambda_{B}B_{p,u}}|+_{u}\rangle\right)\left(\prod_{v}e^{\beta A_{v,m}}|\uparrow_{m}\rangle\right)\\ \end{split}$ (81)
@@
$|\rho(\beta\lambda_{A},\beta\lambda_{B})\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}=\mathcal{U}_{u}(\beta\lambda_{B})\mathcal{U}_{m}(\beta\lambda_{A})|+_{u}\rangle|\uparrow_{m}\rangle$ (82)
@@ V.2 SW-SSB phase
Now we consider the regime $\lambda_{A}\ll h\ll\lambda_{B}$ and apply $\lambda_{A}$ perturbatively. At $\lambda_{A}=0$, the Hamiltonian is classical, and has an extensive number of exactly degenerate ground states [^15] The degeneracy is absent if $\{h_{e}\}$ are taken from a Gaussian distribution Fisher and Huse (1987). (which we will discuss in more details in Sec. V.4). However, the $-\lambda_{A}\sum_{V}A_{v}$ term lifts the exact degeneracy, for general $\{h_{e}\}$ configurations, at extensively large order in degenerate perturbation theory. We use the resulting set of unique ground states to form the density matrix $\rho(\{h_{e}\})$. In this section, we will treat $-\lambda_{A}\sum A_{v}$ as a perturbation that only appears at the level of degenerate perturbation theory, and argue that the perturbed density matrix possesses similar order and disorder parameters as the ones in the random vertex example (42) in the SW-SSB phase. Therefore, the density matrix must be in the SW-SSB phase of the Wilson 1-form symmetry as well.
@@
    Minimize all of the field terms $-\sum h_{e}Z_{e}$, i.e. $Z_{e}=\mathrm{sgn}(h_{e})$.
@@
(${n+m\choose n}={n+m\choose m}$ (94)
different paths of the same shortest length $n+m$. For this configuration of $\{h_{e}\}$, there is therefore a (${n+m\choose n}$-fold ground state degeneracy at $\lambda_{A}=0$. For example, if $n=m=1$, then there is a two-fold degeneracy of classical ground states $|\psi_{1}\rangle,|\psi_{2}\rangle$. A small $-\lambda_{A}\sum A_{v}$ term lifts this degeneracy at first order in ${\lambda_{A}/h}$, resulting in a unique ground state $\frac{1}{\sqrt{2}}(|\psi_{1}\rangle+|\psi_{2}\rangle)$ in a finite-sized system. Since a particular vertex term $A_{v}$ toggles between $|\psi_{1}\rangle$ and $|\psi_{2}\rangle$, the expectation value of this particular vertex term is 1 in this perturbed ground state. More generally, the ground state will be an equal weight superposition of all the (${n+m\choose n}$ degenerate ground states between these two plaquettes with $B_{p}=-1$.
Now consider the entire ground state |$\ket{\psi(\{h_{e}\})}$ of a typical disorder realization $\{h_{e}\}$. The argument above shows that the ground state consists of many short strings to minimally correct the plaquette violations which leads to local coherence of $A_{v}$. In such a state, in order to have $\bra{\psi(\{h_{e}\})}\prod_{v\in R}A_{v}\ket{\psi(\{h_{e}\})}\neq 0$ in a region $R$, we need to have two $B_{p}=-1$ plaquettes on two diagonal corners of $R$ and no $B_{p}=-1$ plaquettes in any part of the region which will cut off the range of coherence of $A_{v}$. Since $\mathrm{sgn}\left(\prod_{e\in p}h_{e}\right)$ has probability $1/2$ of being either $1$ or $-1$ for every plaquette $p$, when averaging over all possible disorder realization, such a probability roughly decays as $\left(\frac{1}{2}\right)^{|R|}$. [^16] There is also a subleading $\left(\frac{1}{2}\right)^{|\partial R|}$ term from the fact that even in such a configuration of $\{h_{e}\}$, $\bra{\psi(\{h_{e}\})}\prod_{v\in R}A_{v}\ket{\psi(\{h_{e}\})}\sim\left(\frac{1}{2}\right)^{2(n+m)}$ according to (94). One can see that the scaling of the order and disorder parameters match with those discussed in Sec. IV.2. In particular, $O_{1,x}\sim\left(\frac{1}{2}\right)^{|R|}$.
@@ V.4 Gaussian distribution
$\begin{split}&O_{1,x}=\tanh^{r^{2}}(\beta_{A})\qquad O_{1,z}=e^{-\mu r/\xi}\\ &O_{2,x}=O_{2,z}=1\\ &D_{1,x}\sim e^{-\mu\mathrm{supp}(D_{1})/\xi}\qquad D_{1,z}\sim e^{-\mu|i-j|/\xi}\\ &D_{2,x}=e^{-\mu|i-j|/\xi}\qquad D_{2,z}=e^{-4\beta_{A}}\\ \end{split}$ (103)
@@ V.5 Relation to Ising spin glass via gauging
[…]
@@
$\tilde{O}_{2}=\frac{\mathrm{Tr}\left(\tilde{\rho}\sigma^{z}_{i}\sigma^{z}_{j}\tilde{\rho}\sigma^{z}_{i}\sigma^{z}_{j}\right)}{\mathrm{Tr}\left(\tilde{\rho}^{2}\right)}=\frac{\sum_{n}p_{n}^{2}\left|\bra{\psi_{n}}\sigma^{z}_{i}\sigma^{z}_{j}\ket{\psi_{n}}\right|^{2}}{\sum_{n}p_{n}^{2}},$ (108)
@@ VI Discussion
For simplicity, we considered just two particular kinds of disorder (random $A_{v}$ term and random $Z_{e}$ field). However, the perturbative methods introduced in Sec. V based on averaging over states obtained by quasi-adiabatic continuation should generalize broadly to other kinds of disorder. For example, we expect that we can apply similar methods to deal with Hamiltonians that also have a small $-\sum_{e}h_{e}X_{e}$ term. This term explicitly breaks the exact strong Wilson symmetry, but we expect to observe an emergent strong Wilson symmetry for sufficiently small disorder amplitude.
@@
where $W_{e}$ is a short string operator for the semion (we can also consider the doubled semion Hamiltonian with random $h_{e}W_{e}$ terms). At $p=\frac{1}{2}$, we actually do not get a SW-SSB/intrinsically mixed phase, like we do when $W_{e}$ is a short string operator $Z_{e}$ in toric code. This can be seen by examining the Choi state. The Choi state for the initial pure doubled semion state describes two copies of doubled semion topological order: $|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}=|ds\rangle|ds\rangle$. The above channel at $p=\frac{1}{2}$ condenses $s_{u}\bar{s}_{l}$ in the Choi state, where $s_{u}$ is the semion of the upper layer and $\bar{s}_{l}$ is the antisemion of the lower layer. Therefore, the state after $p=\frac{1}{2}$ decoherence is symmetric under the $s_{u}\bar{s}_{l}$ weak 1-form symmetry. The remaining strong symmetries are string operators for $\bar{s}_{u}$ and $s_{l}$, which demonstrate ST-SSB (because the string operator of $\bar{s}_{1}s_{2}$ is also SSB). Therefore, we just have ST-SSB rather than SW-SSB, and a WS 1-form symmetry rather than one that is WT-SSB. It is easy to see that the Choi state describes a single copy of doubled semion at $p=\frac{1}{2}$; the mixed density matrix is equivalent to that of a chiral semion topological order, as discussed in Ref. Sohal and Prem (2024).
@@ Acknowledgements
We thank Yu-Hsueh Chen, Meng Cheng, Tarun Grover, Timothy Hsieh, Chao-Ming Jian, Curt von Keyserlingk, Ruochen Ma, Abhinav Prem, Tomohiro Soejima, Senthil Todadri, Ruben Verresen, Ashvin Vishwanath, Chong Wang for inspiring discussions. We especially thank Rahul Sahay for calculations related to Appendix IV.4 and Shengqi Sang for many helpful discussions related to Markov length. Zhu-Xi thanks Jong Yeon Lee and Subir Sachdev for discussion on the replica trick. YX thanks Chao-Ming Jian for collaborating on related works. This research is partially supported by the Simons Collaboration on Ultra Quantum Matter which is a grant from the Simons Foundation (651440, Z.-X. L.), and grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP). YX acknowledges support from the NSF through OAC-2118310. JHZ is supported by the U.S. Department of Energy under the Award Number DE-SC0024324. C.X. is supported by the Simons Foundation through the Simons Investigator program.

### v3

@@ IV.4 Path of gapped parent Hamiltonians
To further support the results in the previous subsection, we show below that there is an explicit path of gapped, local Hamiltonians connecting a parent Hamiltonian for $|\rho(\beta_{A})\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}$ to a parent Hamiltonian for $|\rho_{B}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}$. This clearly demonstrates that these two states are in the same phase, and share the same TEE (modulo spurious contributions). The idea is to find an exact unitary operator $\mathcal{U}(\beta_{A})_{m}$ which has the same action on $|\uparrow_{m}\rangle$ as $\prod e^{\beta_{A}A_{v,m}}$ (upon proper renormalization):
$|\rho(\beta_{A})\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}=\mathcal{U}_{m}(\beta_{A})|tc_{u}\rangle|\uparrow_{m}\rangle$ (67)
@@
$H(\beta)$ is a local parent Hamiltonian for $|\rho(\beta_{A})\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}$, whose terms are exponentially localized with a length scale set by $\beta_{A}$.
@@
Since the circuit (LABEL:seqcirc) is only sequential in the $y$ direction only, it leads to a finite correlation length only in the $y$ direction. In other words, $\mathcal{U}_{m}(\beta_{A})O_{e,m}\mathcal{U}_{m}(\beta_{A})^{\dagger}$ generically maps a local operator $O_{e,m}$ supported on site $e$ to an operator with strict locality in the $x$ direction but an exponential tail in the +y+y direction (note that there is no tail in the y-y direction). In Appendix E, we show that $\mathcal{U}(\beta_{A})_{m}$ is locality-preserving for all local operators. In particular, for a verticle link $e$,
@@
is a gapped Hamiltonian that consists of exponentially local terms, as long as $\beta_{A}$ is finite. When $\beta_{A}=0$, the ground state of $H(\beta_{A})$ is simply a toric code fixed point state stacked with a trivial paramagnet, so it clearly has TEE $\log 2$. The same must hold for the ground state $|\rho(\beta_{A})\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}$ of $H(\beta_{A})$ for any finite $\beta_{A}$.
@@ IV.5 Random plaquette term
$|\rho_{\beta}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}\propto\prod_{v}e^{\beta\lambda_{A}(A_{v,u}+A_{v,l})/2}\prod_{p}e^{\beta\lambda_{B}(B_{p,u}+B_{p,l})/2}|\rho_{1}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}},$ (78)
where $|\rho_{1}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}$ is the Choi state of the identity matrix, which is a state where every spin in $\mathcal{H}_{u}$ is paired with a spin in $\mathcal{H}_{l}$ in a Bell pair. We can write $|\rho_{1}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}$ as the state obtained from full decoherence of toric code with both the $X$ and $Z$ channels.
$|\rho_{1}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}\propto\prod_{e,e^{\prime}}\left(\frac{1+Z_{e,u}Z_{e,l}}{2}\right)\left(\frac{1+X_{e^{\prime},u}X_{e^{\prime},l}}{2}\right)|tc_{u}\rangle|tc_{l}\rangle.$ (79)
As in Sec. IV.3, we will rewrite the state with an additional set of spins acted on by $X_{m},Z_{m}$. Following (61) and (62), $|\rho_{1}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}$ can be written as
$|\rho_{1}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}=|+_{u}\rangle|\uparrow_{m}\rangle$ (80)
Therefore, $|\rho_{\beta}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}$ can be written as
[…]
@@
$|\rho(\beta\lambda_{A},\beta\lambda_{B})\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}=\mathcal{U}_{u}(\beta\lambda_{B})\mathcal{U}_{m}(\beta\lambda_{A})|+_{u}\rangle|\uparrow_{m}\rangle$ (82)
@@ V.2 SW-SSB phase
Now we consider the regime $\lambda_{A}\ll h\ll\lambda_{B}$ and apply $\lambda_{A}$ perturbatively. At $\lambda_{A}=0$, the Hamiltonian is classical, and has an extensive number of exactly degenerate ground states [^15] The degeneracy is absent if $\{h_{e}\}$ are taken from a Gaussian distribution Fisher and Huse (1987). (which we will discuss in more details in Sec. V.4). However, the $\lambda_{A}\sum_{V}A_{v}$ term lifts the exact degeneracy, for general $\{h_{e}\}$ configurations, at extensively large order in degenerate perturbation theory. We use the resulting set of unique ground states to form the density matrix $\rho(\{h_{e}\})$. In this section, we will treat $\lambda_{A}\sum A_{v}$ as a perturbation that only appears at the level of degenerate perturbation theory, and argue that the perturbed density matrix possesses similar order and disorder parameters as the ones in the random vertex example (42) in the SW-SSB phase. Therefore, the density matrix must be in the SW-SSB phase of the Wilson 1-form symmetry as well.
@@
    Minimize all of the field terms $\sum h_{e}Z_{e}$, i.e. $Z_{e}=\mathrm{sgn}(h_{e})$.
@@
${n+m\choose n}={n+m\choose m}$ (94)
different paths of the same shortest length $n+m$. For this configuration of $\{h_{e}\}$, there is therefore a $n+m\choose n$}-fold ground state degeneracy at $\lambda_{A}=0$. For example, if $n=m=1$, then there is a two-fold degeneracy of classical ground states $|\psi_{1}\rangle,|\psi_{2}\rangle$. A small $\lambda_{A}\sum A_{v}$ term lifts this degeneracy at first order in $\lambda_{A}/h$}, resulting in a unique ground state $\frac{1}{\sqrt{2}}(|\psi_{1}\rangle+|\psi_{2}\rangle)$ in a finite-sized system. Since a particular vertex term $A_{v}$ toggles between $|\psi_{1}\rangle$ and $|\psi_{2}\rangle$, the expectation value of this particular vertex term is 1 in this perturbed ground state. More generally, the ground state will be an equal weight superposition of all the $n+m\choose n$} degenerate ground states between these two plaquettes with $B_{p}=-1$.
[…]
@@ V.4 Gaussian distribution
[…]
@@ V.5 Relation to Ising spin glass via gauging
The toric code with random field is related to the Ising model with random coupling via gauging. Specifically, the disordered Hamiltonian in Eq. 86 at $\lambda_{B}\gg\Delta_{h}$ can be mapped to a 2DD random-bond Ising model in transverse field Tsomokos et al. (2011):
@@
[…]
@@ VI Discussion
For simplicity, we considered just two particular kinds of disorder (random $A_{v}$ term and random $Z_{e}$ field). However, the perturbative methods introduced in Sec. V based on averaging over states obtained by quasi-adiabatic continuation should generalize broadly to other kinds of disorder. For example, we expect that we can apply similar methods to deal with Hamiltonians that also have a small $\sum_{e}h_{e}X_{e}$ term. This term explicitly breaks the exact strong Wilson symmetry, but we expect to observe an emergent strong Wilson symmetry for sufficiently small disorder amplitude.
@@
where $W_{e}$ is a short string operator for the semion (we can also consider the doubled semion Hamiltonian with random $h_{e}W_{e}$ terms). At $p=\frac{1}{2}$, we actually do not get a SW-SSB/intrinsically mixed phase, like we do when $W_{e}$ is a short string operator $Z_{e}$ in toric code. This can be seen by examining the Choi state. The Choi state for the initial pure doubled semion state describes two copies of doubled semion topological order: $|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}=|ds\rangle|ds\rangle$. The above channel at $p=\frac{1}{2}$ condenses $s_{u}\bar{s}_{l}$ in the Choi state, where $s_{u}$ is the semion of the upper layer and $\bar{s}_{l}$ is the antisemion of the lower layer. Therefore, the state after $p=\frac{1}{2}$ decoherence is symmetric under the $s_{u}\bar{s}_{l}$ weak 1-form symmetry. The remaining strong symmetries are string operators for $\bar{s}_{u}$ and $s_{l}$, which demonstrate ST-SSB (because the string operator of $\bar{s}_{1}s_{2}$ is also SSB). Therefore, we just have ST-SSB rather than SW-SSB, and a WS 1-form symmetry rather than one that is WT-SSB. It is easy to see that the Choi state describes a single copy of doubled semion at $p=\frac{1}{2}$; the mixed density matrix is equivalent to that of a chiral semion topological order, as discussed in Ref. Sohal and Prem (2024).
@@ Acknowledgements
We thank Yu-Hsueh Chen, Meng Cheng, Tarun Grover, Timothy Hsieh, Chao-Ming Jian, Curt von Keyserlingk, Shang Liu, Ruochen Ma, Abhinav Prem, Tomohiro Soejima, Senthil Todadri, Ruben Verresen, Ashvin Vishwanath, and Chong Wang for inspiring discussions. We especially thank Rahul Sahay for calculations related to Appendix IV.4 and Shengqi Sang for many helpful discussions related to Markov length. Zhu-Xi thanks Jong Yeon Lee and Subir Sachdev for discussion on the replica trick. YX thanks Chao-Ming Jian for collaborating on related works. This research is partially supported by the Simons Collaboration on Ultra Quantum Matter which is a grant from the Simons Foundation (651440, Z.-X. L.), and grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP). YX acknowledges support from the NSF through OAC-2118310. JHZ is supported by the U.S. Department of Energy under the Award Number DE-SC0024324. C.X. is supported by the Simons Foundation through the Simons Investigator program.